# Assessing the Political Fairness of Multilingual LLMs: A Case Study based on a 21-way Multiparallel EuroParl Dataset

### v1

@@ Abstract
The political biases of Large Language Models (LLMs) are usually assessed by simulating their answers to English surveys. In this work, we propose an alternative framing of political biases, relying on principles of fairness in multilingual translation. We systematically compare the translation quality of speeches in the European Parliament (EP), observing systematic differences with majority parties from left, center, and right being better translated than outsider parties. This study is made possible by a new, 21-way multiparallel version of EuroParl, the parliamentary proceedings of the EP, which includes the political affiliations of each speaker. The dataset consists of 1.5M sentences for a total of 40M words and 249M characters. It covers three years, 1000+ speakers, 7 countries, 12 EU parties, 25 EU committees, and hundreds of national parties.
@@ 1.  Introduction
To showcase the potential of 21-EuroParl, we conduct a study on the political fairness of multilingual Large Language Models (LLMs). This question is critical as LLMs are used daily through chatbots by hundreds of millions of users Milmo et al. (2023); Top Websites Ranking (2025) and are increasingly embedded in democratic procedures and processes Small et al. (2023); Tessler et al. (2024); Revel and Penigaud (2025).However, their biases have so far mostly been assessed through simulated surveys, written in English (Feng et al., 2023; Rozado, 2023; Hartmann et al., 2023; Santurkar et al., 2023; Durmus et al., 2024; Motoki et al., 2024; Potter et al., 2024; Boelaert et al., 2025; Röttger et al., 2024; Ceron et al., 2024). The propose an alternative framing of political biases and ask the following question: does LLM translation quality systematically vary depending on the political content of the source text? The answer is not straightforward as existing Machine Translation metrics are not well calibrated across language pairs. We therefore introduce a novel way to aggregate translation scores across multiple language pairs and partisan speeches based on Borda counts. We find systematic differences in scores depending on the source languages, as well as the party affiliation, across several model sizes and families. In summary, our contributions are threefold:
@@ 2.1.  From Parliamentary Speeches into Multilingual Corpora
To the best of our knowledge, this work is the first to leverage Machine Translation as a way to question the political biases of multilingual LLMs. Indeed, most existing works on political biases: (a) focus on monolingual, English-centric LLMs; (b) rely on questionnaires “administered” to LLMs (e.g., querying a model about a degree of agreement with statements such as “Sex outside marriage is usually immoral”).
More precisely, we refer to the following works: (Feng et al., 2023; Rozado, 2023; Hartmann et al., 2023; Santurkar et al., 2023; Durmus et al., 2024; Motoki et al., 2024; Potter et al., 2024). Following works have demonstrated the brittleness of the questionnaire approach, as the computation of LLMs’ answers is sensitive to details of the prompt format (e.g., using a new line instead of a space), also yielding a lack of overall coherence (Boelaert et al., 2025; Röttger et al., 2024; Ceron et al., 2024). Röttger et al. (2025) differ from these works as they assess the stance of LLMs when prompted to generate texts about debated issues, while still limited to English and a US-centric context. We depart from these works and, instead of measuring political biases of LLMs based on the detected political leaning of the text they produce, we frame it in terms of a group-fairness analysis, based on a practical application (Machine Translation), and assess biases over a large set of 420 language pairs. This allows us to also cover a much more diverse set of texts, reflecting the fact that the political spectrum of the European Union is arguably more varied than that of the bipartisan US, which largely maps to a liberal-conservative axis (Ramaciotti et al., 2024; Vendeville et al., 2025). In the EU, other ideological dimensions emerge, e.g., Green-Alternative-Libertarian versus Traditional-Authoritarian-Nationalist (aka GAL-TAN), Urban-Rural, or Pro-Anti Immigration (Jolly et al., 2022).
@@ 3.2.  Sentence Alignment
Table 4: sBLEU scores for Qwen3-8B: source languages in rows and target languages in columns. Scores are comparable within the same column (target language) but not across columns (within the same row, i.e. source language).
@@ 4.1.  Method
Therefore, to measure political fairness based on sBLEU and COMET, we compute a Borda Count of EU parties per language pair (Section 5.3). Borda Count is a simple, yet effective, way to aggregate rankings. For a given model, metric, and language pair, all EU parties are ranked according to the metric score. The EU party with the lowest score gets 0 points, the second-to-last gets 1 point, etc. until the best one gets 7 points (as there are 8 parties). While initially proposed as a voting system, Borda Counts has been used in Information Retrieval to aggregate rankings Aslam and Montague (2001); Bassani and Romelli (2022). It has also been used more recently for benchmarking models across tasks with different metrics (Colombo et al., 2022) or meta-evaluating metrics across language pairs (Freitag et al., 2023). Our method differs from the usual fairness metrics, e.g. demographic parity or group calibration, that are designed for classification settings (Czarnowska et al., 2021; Barocas et al., 2023).
@@
4.2.  Implementation
We experiment with two families of posttrained LLMs: Qwen3 (Yang et al., 2025), in two sizes (4B and 8B), and Llama-3.1-8B (Grattafiori et al., 2024).
We originally carried out experiments with the smaller Llama-3.2 (in sizes 1B and 3B) but found the translation scores overall too poor to be reliably compared across parties (for most language pairs, sBLEU scores were below 10, see Appendix A.1). Llama-3.1 is a multilingual model trained with a very large share of English texts, accounting for 92% of the pretraining data (or 75% when excluding code) and 97% English posttraining data (or 82% when excluding code). Little is known about Qwen3’s data mix, except that it covers 119 languages.
[…]
Approximately 1,500 V100 GPU hours were used for the experiments.
Figure 5: sBLEU against COMET scores of the 420 language pairs for each model. A few language pairs are annotated (source $\rightarrow{}$ target language) along the quantiles.
Model Ranking $p<0.01$ ALDE ECR EFA EFD EPP NA NGL S&D
Fair Model 0 3.50 3.50 3.50 3.50 3.50 3.50 3.50 3.50
Qwen3-4B sBLEU 360 4.98 3.23 3.70 3.06 5.17 1.45 1.41 4.99
Qwen3-8B sBLEU 363 5.13 3.02 3.55 3.08 5.17 1.65 1.18 5.22
Llama-3.1-8B sBLEU 304 4.83 3.73 3.63 2.84 4.56 2.20 1.08 5.11
Qwen3-4B COMET 418 4.51 3.20 1.63 4.49 6.72 0.79 0.94 5.72
Qwen3-8B COMET 419 4.82 3.41 1.47 4.02 6.64 1.03 0.80 5.82
Llama-3.1-8B COMET 419 4.94 4.15 1.13 3.58 6.37 1.72 0.50 5.61
Table 6: Borda Count of political parties rankings averaged over the 420 language pairs (from 0-7, the higher the better), based on sBLEU or COMET rankings. The $p<0.01$ column shows how many language pairs (out of 420) have their performance variance significantly explained by the party variable according to a Kruskal-Wallis test.
@@ 5.1.  Overall Translation Quality
For the 23,386 examples of the test set, we compute sBLEU and COMET scores for the 420 language pairs, giving 9,822,120 observations for each metric. We report the average sBLEU scores of Qwen3-8B per language direction (source languages in rows and target languages in columns) in Table 4 (results for all models are reported in Appendix A). This model overall achieves good translation performance, especially for high-resource pairs such as en-es where sBLEU scores exceed $40$. However, results are not uniform and even for the most favorable case (translating into English), sBLEU scores vary by a wide margin. This is also reflected in Figure 5 where we plot sBLEU against COMET scores for the 420 language pairs for each model. We can see that Qwen3-8B slightly outperforms Llama-3.1-8B and that Qwen3-4B underperforms, but the tendencies across language pairs are largely similar: in fact, the sBLEU scores of Qwen3-4B and Qwen3-8B have a Pearson $\rho$ correlation of .98, as well as that of Qwen3-8B and Llama-3.1-8B. We can also see a strong correlation of the two metrics: the COMET and sBLEU scores of Qwen3-4B have a Pearson $\rho$ correlation of .91, although this number goes down to .77 for Qwen3-8B and Llama-3.1-8B. See (Kocmi et al., 2024) on how to interpret COMET scores.
@@ 5.2.  Language Fairness
Before diving into the political fairness of LLMs, we first study their language fairness, trying to evaluate which language do they translate the worst, and which one do they translate the best. A hasty conclusion would rely on the average of sBLEU/COMET scores across language pairs, to find which language gets the highest score when translating into/out of it. However, COMET scores are not well calibrated across language pairs (Section 4.1). sBLEU, while not well calibrated across target languages, still enables to compare multiple source languages translating towards the same target language. [^11] Assuming that the source texts are similar, as is the case in this section. Therefore, we apply the Borda method described above as follows: for each target language, we rank all source languages and assign 0 points to the worst, up to 19 points to the best. For example in the bg column of Table 4, et-bg has the worst sBLEU score of 17.8, so it gets 0 points, while en-bg has the best score of 29.3 so it gets 19 points.
Results are reported in Table 5. As expected, we find that en is the best translated language, [^12] The only target languages where it is not ranked first, across all models, are cs and sk, where it is then ranked second (and sk – resp. cs – is ranked first, the two languages being closely related). consistently across models, followed by es. Unlike what Table 4 would suggest (the et column has the lowest sBLEU scores), we find that lt is the worst translated language across models, especially for Qwen3-8B where it is almost always ranked last. et is also among the worst translated languages. A surprising result is that de and fr, which have a special status within the EP, have a lower rank than sp, but also than bg and ro, reflecting perhaps a small representation of these two languages in the data mix of Llama and Qwen. We also observe, for all models, that all of the 21 target languages have their performance variance significantly explained by the source language variable according to a Kruskal-Wallis test (a non-parametric version of ANOVA; Kruskal and Wallis, 1952) – confirming that some source languages are more difficult to process than others Bugliarello et al. (2020).
@@ 5.3.  Political Fairness
We now turn to the political fairness of LLMs. We apply the same Borda method to rank political parties across language pairs. This can also be done with COMET, as COMET scores are comparable within the same language pair.
@@
Results are reported in Table 6. We find that EPP, S&D, and ALDE get the best translation quality and NA and NGL the worst, across all models and metrics. The preferences are very marked, with EPP nearing 7 points (i.e. always ranked first) and NGL nearing 0 points (i.e. always ranked last), especially with COMET. The preferences are especially consistent within the same model family: the Borda Count of Qwen3-4B and Qwen3-8B have a Pearson $\rho$ correlation of .99, for both sBLEU- and COMET-based scores. Even across model families, results are very consistent, e.g. the Borda Count based on sBLEU of Qwen3-8B and Llama-3.1-8B have a Pearson $\rho$ correlation of .96, up to .98 for COMET-based scores. For most language pairs, the party variable significantly explains the variance of performance according to a Kruskal-Wallis test ($p<0.01$), especially with COMET where only et-nl, fi-et, and fi-lv are not significative, across all models. An ideally fair model, on the contrary, would not have its performance variance explained by the party variable. It would have near identical performance for all parties, i.e. random rankings for each language pair, which would result in the same Borda Count for all parties: 3.5.
tgt ALDE ECR EFA EFD EPP NA NGL S&D
bg 20.83 21.56 21.28 21.59 21.81 19.80 20.88 21.93
cs 17.21 16.90 16.80 17.29 17.37 16.47 15.99 17.46
da 20.91 20.35 20.60 19.88 20.55 19.27 19.93 20.52
de 21.35 21.73 20.38 21.00 21.10 18.10 19.70 21.34
el 18.70 18.19 18.21 18.16 18.89 17.29 17.27 19.08
en 36.04 35.65 34.67 36.00 36.35 34.65 35.01 36.60
es 30.81 30.05 29.85 31.39 32.20 30.22 30.66 31.52
et 8.62 8.27 8.67 7.91 8.16 8.14 7.66 8.19
fi 12.01 11.42 11.73 10.92 11.45 10.77 10.80 11.47
fr 25.99 25.45 25.35 26.27 27.19 24.74 25.68 26.78
hu 12.39 11.73 12.82 11.97 12.37 12.19 11.55 12.19
it 21.27 20.15 20.70 21.23 21.78 19.94 20.77 21.50
lt 11.86 11.10 11.77 11.40 11.41 11.26 10.55 11.43
lv 11.91 11.09 11.88 11.63 11.87 11.24 10.75 11.59
nl 20.03 19.54 19.05 19.03 19.34 18.89 18.41 19.79
pl 16.12 15.39 15.79 15.59 16.20 15.76 15.22 16.21
pt 24.34 23.47 23.64 24.41 24.60 23.76 22.26 24.55
ro 23.91 23.07 23.58 23.40 23.86 23.09 22.81 24.13
sk 17.03 16.99 16.46 16.50 17.00 16.15 16.08 17.07
sl 14.57 14.30 14.28 13.93 14.64 13.88 13.96 14.50
sv 19.85 19.57 18.89 18.83 19.18 18.43 18.43 19.27
@@
While Table 6 reveals consistent political bias towards some parties across language pairs, we also report the per-target-language sBLEU scores of Qwen3-8B for every EU party in Table 7 (see other models in Appendix B). For every model, and target language, the performance variance is significantly explained by the party variable according to a Kruskal-Wallis test ($p<0.01$). Although multilingual LLMs are very English-centric, we find that their political bias follows similar tendencies across all target languages, e.g. S&D, EPP, and ALDE are often the best translated parties, and NGL and NA often the worst (which is reflected in their Borda Count in Table 6). Computing a Borda ranking from Table 7 gives very close results to Table 6, [^13] For example, in the xx-bg row of Table 7, S&D gets the highest sBLEU score of 21.93 (averaged over all source languages: cs-bg, da-bg, etc.) so it gets 7 points, and NA gets the worst score of 19.80, so it gets 0 points. i.e. our results are consistent regardless if we aggregate sBLEU scores per language pair or only per target language.

### v2

@@ Abstract
The political biases of Large Language Models (LLMs) are usually assessed by simulating their answers to English surveys. In this work, we propose an alternative framing of political biases, relying on principles of fairness in multilingual translation. We systematically compare the translation quality of speeches in the European Parliament (EP), observing systematic differences with majority parties from left and right being better translated than outsider parties. This study is made possible by a new, 21-way multiparallel version of EuroParl, the parliamentary proceedings of the EP, which includes the political affiliations of each speaker. The dataset consists of 1.5M sentences for a total of 40M words and 249M characters. It covers three years, 1000+ speakers, 7 countries, 12 EU parties, 25 EU committees, and hundreds of national parties.
@@ 1.  Introduction
To showcase the potential of 21-EuroParl, we conduct a study on the political fairness of multilingual Large Language Models (LLMs). This question is critical as LLMs are used daily through chatbots by hundreds of millions of users Milmo (2023); Top Websites Ranking (2025) and are increasingly embedded in democratic procedures and processes Small et al. (2023); Tessler et al. (2024); Revel and Penigaud (2025). However, their biases have so far mostly been assessed through simulated surveys, written in English (Feng et al., 2023; Rozado, 2023; Hartmann et al., 2023; Santurkar et al., 2023; Durmus et al., 2024; Motoki et al., 2024; Potter et al., 2024; Boelaert et al., 2025; Röttger et al., 2024; Ceron et al., 2024). We propose an alternative framing of political biases and ask the following question: does LLM translation quality systematically vary depending on the political content of the source text? The answer is not straightforward as existing Machine Translation metrics are not well calibrated across language pairs. We therefore introduce a novel way to aggregate translation scores across multiple language pairs and partisan speeches based on Borda counts. We find systematic differences in scores depending on the source languages, as well as the party affiliation, across several model sizes and families. In summary, our contributions are threefold:
@@ 2.1.  From Parliamentary Speeches into Multilingual Corpora
Finally, Yang et al. (2023) propose Multi-EuP to assess the multilingual fairness of Information Retrieval methods, for which they consider topics of the EP as queries and relevant documents as any speech from this topic.
@@
English surveys
To the best of our knowledge, this work is the first to leverage Machine Translation as a way to question the political biases of multilingual LLMs. Indeed, most existing works on political biases: (a) focus on monolingual, English-centric LLMs; (b) rely on surveys “administered” to LLMs (e.g., querying a model about a degree of agreement with statements such as “Sex outside marriage is usually immoral”). More precisely, we refer to the following works: (Feng et al., 2023; Rozado, 2023; Santurkar et al., 2023; Motoki et al., 2024; Potter et al., 2024).
Brittleness
Following works have demonstrated the brittleness of the survey approach, as the computation of LLMs’ answers is sensitive to the exact phrasing of the prompt, also yielding a lack of overall coherence (Röttger et al., 2024; Ceron et al., 2024; Boelaert et al., 2025).
Multilingual
Durmus et al. (2024); Helwe et al. (2025) rely on MT to translate English surveys in multiple languages. However, this poses a confound since MT may (i) obviously contain errors, including stance-changing errors (Shafiabadi and Yvon, 2026); (ii) be biased towards a political party, as we will find out. Nevertheless, Durmus et al. (2024) find that surveying LLMs in English, Russian, Chinese, or Turkish always leads to occidental responses, i.e. the answers are consistent across languages. On the contrary, Helwe et al. (2025) finds that “language has a strong influence”, although they do not control for the spurious prompt effects described above. Labat et al. (2026) have jointly assessed the variation caused by language and prompt phrasing. They found that the effect of language was question-dependent, i.e., for a subset of questions the answer is consistent across languages while for another subset the answer differs across languages.
Other Methods
In addition to their survey-based method, Potter et al. (2024) conduct a human-computer interaction experiment. They find that after debating with an LLM in English, Trump supporters increased their Biden-leaning. Ceron et al. (2025) study the political content of LLMs English pretraining and posttraining data. Using a left-right content classifier, they find a majority of left-leaning data, which concurs with the results discussed above. Röttger et al. (2025) assess the stance of LLMs when prompted to generate English texts about debated issues. They find that LLMs align closely with US Democrats, as opposed to Republicans.
Fairness
We depart from these works and, instead of measuring political biases of LLMs based on the detected political leaning of the text they produce, we frame it in terms of a group-fairness analysis, based on a practical application (Machine Translation), and assess biases over a large set of 420 language pairs. This allows us to also cover a much more diverse set of texts, reflecting the fact that the political spectrum of the European Union is arguably more varied than that of the bipartisan US, which largely maps to a liberal-conservative axis (Ramaciotti et al., 2024; Vendeville et al., 2025). In the EU, other ideological dimensions emerge, e.g., Green-Alternative-Libertarian versus Traditional-Authoritarian-Nationalist (aka GAL-TAN), Urban-Rural, or Pro-Anti Immigration (Jolly et al., 2022).
@@ 3.2.  Sentence Alignment
bg 26.3 28.6 27.9 30.4 44.6 38.6 15.6 17.7 32.3 20.9 28.1 18.0 19.5 26.8 24.3 32.8 33.1 24.6 25.5 28.4
cs 27.8 27.2 27.0 27.6 41.2 36.3 15.4 17.4 30.8 20.5 26.7 16.7 18.1 26.4 23.9 30.7 30.5 31.2 8.2 27.3
da 27.5 24.3 27.7 27.2 44.2 36.9 15.0 17.9 31.6 19.7 27.2 16.0 17.8 26.9 22.2 31.0 30.7 22.6 23.3 32.6
de 25.1 25.0 28.7 26.6 43.2 37.2 15.3 17.7 31.5 20.0 27.1 16.3 17.5 28.6 22.4 30.9 31.1 23.4 23.9 28.5
el 20.2 14.7 22.5 23.4 45.3 37.7 10.5 13.9 31.3 12.7 26.5 10.9 11.8 19.3 15.1 29.4 30.1 16.2 21.2 22.3
en 37.0 32.2 37.5 36.2 38.8 50.1 19.0 22.2 40.4 25.2 34.6 20.4 23.9 32.3 28.7 41.7 44.3 30.3 31.2 37.5
es 27.8 24.6 29.3 21.8 30.8 48.7 15.7 18.1 34.1 20.0 30.0 17.1 18.2 27.2 23.7 36.1 35.4 24.2 24.6 29.3
et 23.9 21.7 23.9 23.7 23.6 34.9 32.5 17.4 27.3 18.9 23.6 15.5 16.7 23.4 20.3 26.9 26.4 20.0 19.9 24.1
fi 23.5 21.4 24.3 23.7 24.0 35.6 32.4 15.1 27.5 18.8 23.8 14.8 16.0 24.0 19.8 27.2 26.4 19.8 20.0 25.5
fr 25.0 23.4 27.1 26.7 28.3 42.8 38.8 13.0 16.2 19.6 29.0 14.9 16.0 26.8 22.2 33.0 32.0 21.8 22.8 27.2
hu 24.1 21.7 24.3 23.8 24.1 36.1 33.2 13.6 16.0 28.1 24.4 15.0 16.1 24.1 20.3 28.4 26.6 19.9 20.5 24.5
it 22.7 21.9 25.4 22.0 25.6 39.9 36.2 13.6 13.5 31.5 18.1 14.3 15.9 25.8 21.1 31.1 29.9 21.0 20.9 25.6
lt 23.0 20.1 22.2 21.7 21.8 33.2 30.7 13.6 14.6 26.0 16.9 22.2 17.1 21.7 19.1 25.5 24.9 18.8 19.3 22.6
lv 24.0 21.7 24.1 23.4 23.6 35.9 32.7 14.9 15.8 27.2 18.5 23.4 17.0 23.1 20.5 27.2 26.6 20.0 20.5 24.0
nl 24.3 21.9 25.4 25.6 24.9 37.4 33.8 13.5 16.1 29.2 18.6 25.5 15.0 16.0 20.5 28.7 27.3 20.2 20.9 25.8
pl 26.2 24.0 25.3 21.7 26.1 38.1 34.6 14.4 16.2 29.4 19.3 25.5 16.2 17.3 24.7 29.4 28.6 22.5 22.7 25.7
pt 29.0 25.2 28.5 26.1 30.8 46.8 41.6 15.2 17.6 34.3 20.4 29.9 16.8 18.1 27.3 23.3 33.9 23.4 23.9 28.4
ro 29.6 25.8 29.6 25.5 32.3 48.7 42.1 15.5 17.5 34.8 21.2 30.0 17.1 19.1 27.7 23.5 35.7 24.3 24.7 29.7
sk 28.4 32.1 27.6 27.8 27.8 42.0 37.0 14.8 17.1 31.2 19.9 26.9 16.6 18.5 26.4 24.3 31.1 31.0 24.9 27.4
sl 27.7 25.7 27.1 27.2 26.9 41.2 36.6 14.5 16.6 30.7 19.4 26.4 16.8 18.0 25.9 23.3 30.4 30.4 24.7 27.0
sv 26.3 23.4 31.4 26.5 26.3 42.1 35.7 14.9 17.4 30.7 19.1 26.6 15.7 17.2 26.3 21.7 30.1 29.7 21.5 22.1
Table 4: sBLEU scores for Llama-3.1-70B: source languages in rows and target languages in columns. Scores are comparable within the same column (target language) but not across columns (within the same row, i.e. source language).
@@ 4.1.  Method
Borda Count
Therefore, to measure political fairness based on sBLEU and COMET, we compute a Borda Count of EU parties per language pair (Section 5.3). Borda Count is a simple, yet effective, way to aggregate rankings (McLean, 2019). For a given model, metric, and language pair, all EU parties are ranked according to the metric score. The EU party with the lowest score gets 0 points, the second-to-last gets 1 point, etc. until the best one gets 7 points (as there are 8 parties). While initially proposed as a voting system, Borda Counts has been used in Information Retrieval to aggregate rankings Aslam and Montague (2001); Bassani and Romelli (2022). It has also been used more recently for benchmarking models across tasks with different metrics (Colombo et al., 2022) or meta-evaluating metrics across language pairs (Freitag et al., 2023). Our method differs from the usual fairness metrics, e.g. demographic parity or group calibration, that are designed for classification settings (Czarnowska et al., 2021; Barocas et al., 2023).
4.2.  Implementation
We experiment with three families of posttrained LLMs: Gemma-3-4B (Team, 2025), Qwen3 (Yang et al., 2025), in two sizes (4B and 8B), and Llama-3.1 (Grattafiori et al., 2024), in two sizes as well (8B and 70B).
We originally carried out experiments with the smaller Llama-3.2 (in sizes 1B and 3B) but found the translation scores overall too poor to be reliably compared across parties (for most language pairs, sBLEU scores were below 10, see Appendix B.1). Llama-3.1 is a multilingual model trained with a very large share of English texts, accounting for 92% of the pretraining data (or 75% when excluding code) and 97% English posttraining data (or 82% when excluding code). Little is known about Gemma or Qwen3’s data mix, except that Qwen3’s covers 119 languages.
[…]
A couple of thousands V100 and H100 GPU hours were used for the experiments.
@@ 5.1.  Overall Translation Quality
For the 23,386 examples of the test set, we compute sBLEU and COMET scores for the 420 language pairs, giving 9,822,120 observations for each metric. We report the average sBLEU scores of Llama-3.1-70B per language direction (source languages in rows and target languages in columns, regardless of the party) in Table 4 (results for all models are reported in Appendix B). This model overall achieves good translation performance, especially for high-resource pairs such as es-en where sBLEU scores exceed $40$. However, results are not uniform and even for the most favorable case (translating into English), sBLEU scores vary by a wide margin. We see that, overall, larger models outperform smaller models (e.g., Llama-3.1-70B vs. -8B, or Qwen3-8B, vs. -4B); Qwen3-8B slightly outperforms Llama-3.1-8B and Gemma-3-4B outperforms Qwen3-4B; but the tendencies across language pairs are largely similar: in fact, the sBLEU scores of Llama-3.1-8B and Llama-3.1-70B have a Pearson $\rho$ correlation of .96. Even across families, the sBLEU scores of Qwen3-8B and Llama-3.1-8B have a Pearson $\rho$ correlation of .98. We can also see a strong correlation of the two metrics: the COMET and sBLEU scores of Qwen3-4B have a Pearson $\rho$ correlation of .91, although this number goes down to .77 for Qwen3-8B and Llama-3.1-8B. See (Kocmi et al., 2024) on how to interpret COMET scores.
@@ 5.2.  Language Fairness
Before diving into the political fairness of LLMs, we first study their language fairness, trying to evaluate which language do they translate the worst, and which one do they translate the best. A hasty conclusion would rely on the average of sBLEU/COMET scores across language pairs, to find which language gets the highest score when translating into/out of it. However, COMET scores are not well calibrated across language pairs (Section 4.1). sBLEU, while not well calibrated across target languages, still enables to compare multiple source languages translating towards the same target language. [^12] Assuming that the source texts are similar, as is the case in this section. Therefore, we apply the Borda method described above as follows: for each target language, we rank all source languages and assign 0 points to the worst, up to 19 points to the best. For example in the bg column of Table 4, el-bg has the worst sBLEU score of 20.2 (regardless of the party), so it gets 0 points, while en-bg has the best score of 37.0 (regardless of the party) so it gets 19 points.
Model Ranking $p<0.01$ NGL S&D EFA ALDE EPP ECR EFD NA
Fair Model 0 3.50 3.50 3.50 3.50 3.50 3.50 3.50 3.50
Gemma-3-4B sBLEU 333 1.09 4.77 3.44 5.00 4.41 4.16 3.27 1.87
Qwen3-4B sBLEU 360 1.41 4.99 3.70 4.98 5.17 3.23 3.06 1.45
Qwen3-8B sBLEU 363 1.18 5.22 3.55 5.13 5.17 3.02 3.08 1.65
Llama-3.1-8B sBLEU 304 1.08 5.11 3.63 4.83 4.56 3.73 2.84 2.20
Llama-3.1-70B sBLEU 312 1.06 5.08 3.81 4.73 4.65 3.33 3.33 2.00
Gemma-3-4B COMET 403 0.57 5.22 1.29 5.26 6.04 4.27 3.72 1.63
Qwen3-4B COMET 418 0.94 5.72 1.63 4.51 6.72 3.20 4.49 0.79
Qwen3-8B COMET 419 0.80 5.82 1.47 4.82 6.64 3.41 4.02 1.03
Llama-3.1-8B COMET 419 0.50 5.61 1.13 4.94 6.37 4.15 3.58 1.72
Llama-3.1-70B COMET 415 0.34 5.47 1.06 5.16 5.81 4.32 3.44 2.40
Table 6: Borda Count of political parties rankings averaged over the 420 language pairs (from 0-7, the higher the better), based on sBLEU or COMET rankings. Columns approximately position political parties from left to right. The $p<0.01$ column shows how many language pairs (out of 420) have their performance variance significantly explained by the party variable according to a Kruskal-Wallis test.
Results are reported in Table 5. As expected, we find that en is the best translated language, [^13] The only target languages where it is not ranked first, across all models, are cs and sk, where it is then ranked second (and sk – resp. cs – is ranked first, the two languages being closely related). consistently across models, followed by es. Unlike what Table 4 would suggest (the et column has the lowest sBLEU scores), we find that lt is the worst translated language across models, especially for Qwen3-8B where it is almost always ranked last. et is also among the worst translated languages. A surprising result is that de and fr, which have a special status within the EP, have a lower rank than es, but also than bg and ro. We also observe, for all models, that all of the 21 target languages have their performance variance significantly explained by the source language variable according to a Kruskal-Wallis test (a non-parametric version of ANOVA; Kruskal and Wallis, 1952) – confirming that some source languages are more difficult to process than others Bugliarello et al. (2020).
@@ 5.3.  Political Fairness
We now turn to the political fairness of LLMs. We apply the same Borda method to rank political parties across language pairs. For a given model, metric, and language pair, all EU parties are ranked according to the metric score (Section 4.1). Therefore, the language fairness (previous section) does not influence political fairness, as each language pair equally contributes to the final Borda count. This can also be done with COMET, as COMET scores are comparable within the same language pair.
@@
Results are reported in Table 6. We do not find a gradual bias from left to right. Rather, we find that EPP, S&D, and ALDE get the best translation quality and NA and NGL the worst, across all models and metrics. The preferences are very marked, with EPP nearing 7 points (i.e. always ranked first) and NGL nearing 0 points (i.e. always ranked last), especially with COMET. The preferences are especially consistent within the same model family: the Borda Count of Llama-3.1-8B and Llama-3.1-70B have a Pearson $\rho$ correlation of .98, for both sBLEU- and COMET-based scores. Even across model families, results are very consistent, e.g. the Borda Count based on sBLEU of Qwen3-8B and Llama-3.1-8B have a Pearson $\rho$ correlation of .96, up to .98 for COMET-based scores. For most language pairs, the party variable significantly explains the variance of performance according to a Kruskal-Wallis test ($p<0.01$), especially with COMET. An ideally fair model, on the contrary, would not have its performance variance explained by the party variable. It would have near identical performance for all parties, i.e. random rankings for each language pair, which would result in the same Borda Count for all parties: 3.5.
tgt NGL S&D EFA ALDE EPP ECR EFD NA
bg 25.66 26.52 26.04 25.63 26.26 26.52 26.19 24.93
cs 22.30 24.18 23.52 23.74 24.01 23.63 23.90 23.39
da 26.17 27.05 27.54 27.69 27.04 27.46 26.42 25.95
de 24.22 25.97 24.93 25.75 25.58 26.01 25.73 22.32
el 25.51 27.79 26.67 27.58 27.74 26.98 26.81 25.55
en 39.76 41.59 39.90 41.12 41.31 40.07 41.54 39.78
es 35.80 36.72 35.02 36.03 37.48 34.94 37.11 35.26
et 14.02 14.47 15.29 15.21 14.63 15.16 14.42 14.72
fi 16.00 16.88 17.65 17.60 16.76 17.05 16.38 16.91
fr 30.38 31.20 29.89 30.18 31.63 29.54 31.07 28.76
hu 18.09 19.48 19.87 19.67 19.51 18.62 19.11 19.49
it 26.44 27.20 26.56 26.61 27.20 25.41 26.57 25.49
lt 15.01 16.28 16.50 16.16 15.96 16.39 15.90 16.25
lv 16.03 17.45 17.94 17.71 17.60 17.05 17.42 17.01
nl 23.98 26.36 25.37 26.62 25.63 25.76 25.39 24.86
pl 21.01 22.28 21.71 21.91 22.28 21.10 21.69 21.38
pt 28.08 31.23 30.64 31.16 31.16 29.69 31.05 30.41
ro 29.80 30.70 30.64 30.51 30.41 30.07 30.47 30.39
sk 21.71 22.73 22.13 22.57 22.66 23.04 22.08 22.06
sl 21.11 22.37 21.76 22.40 22.20 22.05 21.58 20.69
sv 25.84 27.29 27.17 27.79 27.25 27.59 26.94 26.14
Table 7: sBLEU scores of Llama-3.1-70B per target language. Columns approximately position political parties from left to right. Each target language has its performance variance significantly explained by the party variable according to a Kruskal-Wallis test ($p<0.01$).
@@
While Table 6 reveals consistent political bias towards some parties across language pairs, we also report the per-target-language sBLEU scores of Qwen3-8B for every EU party in Table 7 (see other models in Appendix C). For every model, and target language, the performance variance is significantly explained by the party variable according to a Kruskal-Wallis test ($p<0.01$). Although multilingual LLMs are very English-centric, we find that their political bias follows similar tendencies across all target languages, e.g. S&D, EPP, and ALDE are often the best translated parties, and NGL and NA often the worst (which is reflected in their Borda Count in Table 6). Computing a Borda ranking from Table 7 gives very close results to Table 6, [^14] For example, in the xx-cs row of Table 7, S&D gets the highest sBLEU score of 24.18 (averaged over all source languages: cs-bg, da-bg, etc.) so it gets 7 points, and NGL gets the worst score of 22.30, so it gets 0 points. i.e. our results are consistent regardless if we aggregate sBLEU scores per language pair or only per target language.